HUBBLE SPACE TELESCOPE DAILY REPORT #2035 PERIOD COVERED: 0000Z (UTC) 01/07/98 - 0000Z (UTC) 01/08/98 Daily Status Report as of 008/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 Decontamination Procedure. Ultraviolet-blocking contaminants are removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images are taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.3 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/2 7420 (A Near-Infrared Search for Very Low Mass Companions to Stars Within 10 Parsecs of the Sun) The Near Infrared Camera (Camera Two) was used to make searching observations for very low mass companions to variable star LHS168. This star is known to be within 10 parsecs of the Sun. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed NIC/2 7413 (The Dusty Environments of Protostars) The Near infrared Camera (Camera Two) was used to make observational images of the dusty environment around protostar IRAS04381. These observations will address physical conditions in protostellar envelopes and provide deeper insight into star and protoplanetary disk formation. The observations were completed as planned, and no problems were noted. 1.6 Completed NIC/1/2/3 7421 (The Relation Between Quasar Luminosity and Host Galaxy Mass) The Near Infrared Camera (Cameras One, Two, and Three) was used to obtain ground-based, near-infrared imaging of the nearby quasar PG0947+396 which has indicated a relation between the luminosity of its nucleus and the luminous mass of its host galaxy. The observations completed as planned, with no reported problems. 1.7 Completed WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.8 Completed Two Sets of WF/PC-2 7786 (Pure Parallel Archive Survey - Continuation) The WF/PC-2 was used to make observations as part of an archival survey program in which the images were taken in parallel with the prime science observations. These WF/PC-2 parallel observations were obtained in the ultraviolet in a search for interesting objects. The observations were completed as planned, and no problems were reported. 1.9 Completed STIS/MA1/MA2 7657 (STIS Spectroscopic and Imaging Sensitivity, MAMA) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of BD+75D325 in order to study the spectroscopic and imaging sensitivity. Sensitivity measurements are done for all primary tilts of the gratings and the prism used with the MAMAs, except for the G140L and G230L gratings. Imaging sensitivity measurements are done with all apertures, including a neutral density filter. A wavelength and a dark calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Three Sets of NIC/2/3 7922 (New NICMOS Modified One Orbit Parallel) The Near Infrared Camera (Cameras One and Two) was used to make one orbit parallel observations. This proposal obtains "JH" imaging, and G096 and G141 spectroscopy of sky regions when another instrument is prime. STEP64 sequences are used and two dither positions are obtained for each filter. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Two Sets of WF/PC-2 6802 (Cosmology With the Medium Deep Survey) The WF/PC-2 was used to make observations taken in parallel with the prime science observations as part of the Medium Deep Sky Survey. The observations were executed as scheduled, and no problems were reported. 1.12 Completed NIC/2 7498 (NICMOS Imaging Survey of Distant Radio Galaxies. Ellipticals at Z > 2 ?) The Near Infrared Camera (Camera Two) was used to make high resolution and high signal-to-noise ratio imaging survey observations of the radio galaxy USS1410-00. The existence and photometric properties of massive elliptical galaxies at high redshift provide a powerful probe of galaxy evolution. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed STIS/CCD 7783 (STIS Parallel Science) The Space Telescope Imaging Spectrograph (CCD) was used to make observations taken in parallel with the prime science. The observations were completed, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 12 Successful : 12 2.3 Operations Notes: The STIS bad pixel was uplinked at 008/0209Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH